OpenStudy (anonymous):

Suppose the area of a circle is 9TT. What is its diameter?

OpenStudy (chihiroasleaf):

do you mean the area of the circle is \(9 \pi\) ?

OpenStudy (dtan5457):

Area of circle\[\pi r^2\] If area is 9 pi. Radius must be 3. as \[\pi (3)^2=9\pi \]

OpenStudy (dtan5457):

If the diameter is twice as long as it's radius, and the radius is 3, what's the diameter?